OpenStudy (anonymous):

If x > 0 and y >0 and x^9 = 4 and x^7 = 9/y^2, which of the following is an expression for the value of x in terms of y? (a) (4/9)y (b) (2/3)y (c) (3/2)y (d) 6y (e) 36y^2

OpenStudy (anonymous):

\[x^9=x^(7+2)=x^7x^2=4\] On the other hand \[x^9=x^7x^2=(9/y^2)x^2=4\] And the rest should be simple.

OpenStudy (anonymous):

is it (a) 4/9y ?